Ms. Anthony wants to create a special blend using two coffees, one priced ar $8.40 per pound and the other at $7.28 per pound. How many pounds of the $7.28 coffee should she mix with 9 pounds of the $8.40 coffee to sell the mixture for $7.95 per pound?
Answer by checkley77(12844)

8.40*9+7.28x=7.95(9+x)
75.6+7.28x=71.55+7.95x
7.28x-7.95x=71.55-75.60
-.67x=-4.05
x=-4.05/-.67
x=6.045 pounds of $7.28 coffee is used.
Proof:
8.40*9+7.28*6.045=7.95(9+6.045)
75.6+44=7.95*15.045
119.6=119.6